OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point 3x^2-x^3, (1,2)

OpenStudy (anonymous):

take the derivative, \[f'(x)\] then find \[f'(1)\] for the slope and then use the point - slope formula

OpenStudy (anonymous):

Using the power rule to find the derivative. You should get 6x -3x^2.. You are given the point (1,2) so plug in the X value to find the slope of the tangent line which is 3. Use the point-slope formula to obtain the equation of the tangent line.

jimthompson5910 (jim_thompson5910):

First derive y to find the derivative function. This will help you find the slope. y = 3x^2-x^3 y' = 6x-3x^2 y' = -3x^2+6x ---------------- Now find slope of tangent line at x = 1 y' = -3x^2+6x y' = -3(1)^2+6(1) y' = -3(1)+6(1) y' = -3+6 y' = 3 So the slope of the tangent line at x = 1 is 3. So m = 3 ---------------- Now use the equation y = mx+b to find the equation of the line y = mx+b 2 = 3(1)+b 2 = 3+b 2-3 = b -1 = b b = -1 So the equation of the tangent line is y = 3x-1